# An improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$

### v2

@@ II Formalism and perturbative calculations
where $q$ denotes the light down quark $d(s)$ for the CKM-favored (-suppressed) processes.
@@
where the ratios $r_{2}=m_{L}/m_{B_{c}}$ and $r_{3}=m_{\eta_{c}}/m_{B_{c}}$ with $m_{L}(m_{\eta_{c}})$ being the light ($\eta_{c}$) meson mass, $P_{1}$ denotes the momentum carried by $B_{c}$ meson in its rest frame, and $P_{2}$ and $P_{3}$ denotes the momenta carried by $L$ and $\eta_{c}$ mesons moving along the plus and minus $z$-directions, respectively. Thanks to conservation of the angular momentum, the possible polarization vectors $\epsilon_{2L}$ ( Here, the subscript $L$ stands for the longitudinal polarization. Not to be confused with the abbreviation $L$ of light mesons.) could be easily derived through the constraints $P_{2}\cdot\epsilon_{2L}=0$ and $\epsilon_{2L}^{2}=-1$. Notice that, if $L$ is a tensor meson, then a new longitudinal polarization vector could be constructed similarly relative to those of vectors and axial-vectors but with an additional factor $\sqrt{2/3}$ Berger:2000wt ; Datta:2007yk ; Liu:2017cwl ; Liu:2023kxr . Then, specifically, $\epsilon_{2L}=\frac{1}{\sqrt{2(1-r_{3}^{2})}r_{2}}(1-r_{3}^{2},-r_{2}^{2},{\bf 0}_{T})$ for vectors and axial-vectors, while $\epsilon_{2L}=\frac{1}{\sqrt{3(1-r_{3}^{2})}r_{2}}(1-r_{3}^{2},-r_{2}^{2},{\bf 0}_{T})$ for tensors in this work. The momenta of valence quarks in the initial- and final-state mesons are parameterized as
@@
The $B_{c}^{+}\to\eta_{c}L^{+}$ decay amplitude in the iPQCD formalism can therefore be conceptually written as follows,
@@
where $b_{i}$ is the conjugate space coordinate of transverse momentum $k_{iT}$; $t$ is the largest running energy scale in hard kernel $H(x_{i},b_{i},t)$; Tr denotes the trace over Dirac and $SU(3)$ color indices; $C(t)$ stands for the Wilson coefficients including the large logarithms $\ln(m_{W}/t)$ Keum:2000ph ; and $\Phi$ is the wave function describing the hadronization of quark and anti-quark to the meson. The Sudakov factor $e^{-S(t)}$ arises from $k_{T}$ resummation, which provides a strong suppression on the long distance contributions in the small $k_{T}$ (or large $b$) region Botts:1989kf . The detailed discussions for $e^{-S(t)}$ can be easily found in the original Refs. Botts:1989kf ; Liu:2018kuo ; Liu:2020upy . Thus, with Eq. (12), we can give the convoluted amplitudes of the decays $B_{c}^{+}\to\eta_{c}L^{+}$ explicitly through the evaluations of hard kernel $H(x_{i},b_{i},t)$ at leading order in the $\alpha_{s}$ expansion within the iPQCD formalism.
@@
where $P$ is the momentum of $B_{c}$ meson, $N_{c}=3$ is the color factor, and $x$ and ${\bf k}_{T}$ are the momentum fraction and intrinsic transverse momentum of charm quark in the $B_{c}$ meson. Here, $\phi_{B_{c}}(x,{\bf k}_{T})$ is the $B_{c}$-meson leading-twist distribution amplitude, whose explicit form in the impact $\bf b$} space is as the following Liu:2018kuo ,
@@
with the decay constant $f_{B_{c}}=0.489\pm 0.005$ GeV Chiu:2007km , the shape parameter $\beta_{B_{c}}=1.0\pm 0.1$ GeV Liu:2018kuo , and $m_{c}$ and $m_{b}$ the charm and bottom quark masses. Moreover, the normalization constant $N_{B_{c}}$ is fixed by the following relation,
@@
For the $\eta_{c}$ meson, its wave function has been studied within the non-relativistic QCD approach Bondar:2004sv and derived as,
@@
with $P$ and $m$ being the momentum and mass of $\eta_{c}$, and $x$ describing the charm-quark momentum fraction in $\phi^{v}_{\eta_{c}}(x)$ and $\phi^{s}_{\eta_{c}}(x)$ the twist-2 and twist-3 distribution amplitudes,
@@
The light-cone wave functions including distribution amplitudes for light pseudoscalars, scalars, vectors, axial-vectors, and tensors calculated in the QCD sum rules up to twist-3 have been collected in Liu:2023kxr (and references therein). For simplicity, their explicit expressions would no longer be presented in this work. The readers could refer to Liu:2023kxr for detail.
Figure 1: (Color online) Leading order Feynman diagrams for the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism.
The leading-order Feynman diagrams for the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism are shown in Fig. 1. Similar to those in Liu:2023kxr , we use $F_{e}$ and $M_{e}$ to describe the factorizable emission and the nonfactorizable emission amplitudes induced by the (VA)(VA)$(V-A)$ operators. The $B_{c}^{+}\to\eta_{c}L^{+}$ decay amplitude can thus be decomposed into
@@
in which, the related factorization formulas are given explicitly as follows,
@@
    where the ratios $r_{b}=m_{b}/m_{B_{c}}$ and $r_{c}=m_{c}/m_{B_{c}}$. The hard function $h_{i}(x_{i},b_{i})$ and the evolution function $E_{f}(t_{i})$ could refer to those expressions in Ref. Liu:2023kxr .
@@
with the $B_{c}$-meson lifetime $\tau_{B_{c}}$ and the decay width $\Gamma$. Note that, $|{\bf P}_{c}|\equiv|{\bf P}_{L}|=|{\bf P}_{\eta_{c}}|=\sqrt{\lambda(m_{B_{c}}^{2},m_{L}^{2},m_{\eta_{c}}^{2})}/(2m_{B_{c}})$ is the momentum of either $L$ or $\eta_{c}$ meson in the final states, with the $\rm K\ddot{a}ll\acute{e}n$} function $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$ ParticleDataGroup:2024cfk .
@@ III Numerical Results and Discussions
In numerical calculations, central values of the input parameters will be used implicitly unless otherwise stated. We adopt the relevant QCD scale (GeV), masses (GeV), and $B_{c}$-meson lifetime (ps) Keum:2000ph ; ParticleDataGroup:2024cfk
@@
In the following context, we will classify the considered decays into two different groups conventionally, that is, factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$ and factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$, respectively, to present the related iPQCD predictions and phenomenological insights.
III.1 Factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$
Specifically, the factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$ decays include $B_{c}^{+}\to\eta_{c}(P,V)^{+}$, $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$, and $B_{c}^{+}\to\eta_{c}K_{1}(1270,1400)^{+}$, respectively. The decays $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ have been studied extensively with various $B_{c}\to\eta_{c}$ form factors in different approaches, however, achieving different individual BRs and the associated relative ratios as given in Tables 1 and 2. It is clear that the values given in the literature have a wide spread. While, to our best knowledge, the $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ and $\eta_{c}K_{1}(1270,1400)^{+}$ modes have not yet been investigated within the QCD-based factorization framework.
The CP-averaged $B_{c}^{+}\to\eta_{c}(\pi,K)^{+}$ BRs in the iPQCD formalism are presented as,
@@ III.1 Factorizable-emission dominated $B_{c}^{+}\to\eta_{c}L^{+}$
where the uncertainties are dominated by the shape parameter $\beta_{B_{c}}$ from the $B_{c}$-meson distribution amplitude. Although the $B_{c}^{+}\to\eta_{c}\pi^{+}$ BR is not measured yet, its iPQCD result is consistent generally with several predictions from various models and/or approaches already presented in the literature, for example, see Refs. Chang:1992pt ; Ivanov:2006ni ; Rui:2014tpa ; Issadykov:2018myx ; Wu:2024gcq ; Deng:2025znr . Note that, though having a smaller decay constant $f_{\eta_{c}}\sim 0.387$ than $f_{J/\psi}\sim 0.405$ and the same leading-twist distribution amplitudes in both $\eta_{c}$ and $J/\psi$, the $B_{c}^{+}\to\eta_{c}\pi^{+}$ BR is still larger than the $B_{c}^{+}\to J/\psi\pi^{+}$ one in the iPQCD formalism. In fact, from the numerical results of decay amplitudes shown in Table 3, it is found that ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ are governed by highly different QCD dynamics, that is, the former (latter) determined by the contributions from twist-3 (twist-2) distribution amplitude of $\eta_{c}(J/\psi)$ even with enhancement (reduction) by twist-2 (twist-3) one. Any precise measurements in various experiments to help deeply understand the QCD behavior of these $c\bar{c}$-mesons are urgently demanded.
@@
Notice that, the results about BR and relative ratio shown in Eqs. (29) and (31) are simultaneously well consistent with those predicted in Issadykov:2018myx within uncertainties. According to the LHCb data LHCb:2014mvo , that is, $\frac{\sigma(B_{c}^{+})}{\sigma(B^{+})}\times{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})=(6.97\pm 0.13)\times 10^{-6}$ Liu:2025ipe in the fiducial region corresponding to the transverse momentum $0<p_{T}<20$ GeV and the rapidity $2.0<y<4.5$, and the assumption ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})\sim{\cal O}(10^{-3})$, then the promising measurements of ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ in the LHCb experiment might be
@@
Based on the data for $\eta_{c}$ strong decays to stable hadrons ParticleDataGroup:2024cfk , one can derive the multibody $B_{c}$-meson decays via $\eta_{c}$ resonance under the narrow-width approximation.
@@
    lead to the three-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p})$ and five-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-})$ BRs as,
@@
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to p\bar{p}\pi^{+}\pi^{-})=(7.51^{+2.21}_{-1.83})\times 10^{-6}\;,$ (35)
@@
    result in the five-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to 2(\pi^{+}\pi^{-}),(\pi^{+}\pi^{-}K^{+}K^{-}),2(K^{+}K^{-}))$ BRs as,
@@
    $=$ $(2.84^{+1.10}_{-0.99})\times 10^{-6}\;,$
In principle, the related measurements are previously difficult due to the large background and the small efficiency at LHC. However, these BRs around ${\cal O}(10^{-6})$ and above with fully charged final-states are expected to be probed in the near future, since the LHCb detector has accomplished a successful upgrade in 2022.
@@
The $B_{c}^{+}\to\eta_{c}V^{+}$ BRs in the iPQCD formalism can be read as follows,
@@
where the theoretical errors are also dominated mainly by the $B_{c}$-meson shape parameter $\beta_{B_{c}}$. These predictions are well consistent with those in Refs. Chang:1992pt ; Ivanov:2006ni ; Issadykov:2018myx ; Deng:2025znr within uncertainties.
By employing Eqs. (33) and (36), the measurable BRs of multibody $B_{c}^{+}\to\eta_{c}\rho^{+}$ decays through resonance state $\eta_{c}$ could be derived under the narrow-width approximation as follows,
@@
The ratio between the $B_{c}^{+}\to\eta_{c}\rho^{+}$ and $B_{c}^{+}\to(J/\psi,\eta_{c})\pi^{+}$ BRs could be deduced as,
@@
Here, the latter ratio $R_{\rho/\pi}^{\eta_{c}}$ is close to that obtained from the $B_{c}^{+}\to J/\psi(\pi,\rho)^{+}$ decays. Moreover, the ratio between the $B_{c}^{+}\to\eta_{c}K^{*+}$ and $B_{c}^{+}\to\eta_{c}\rho^{+}$ BRs is,

### v3

@@ II Formalism and perturbative calculations
where $q$ denotes the light down quark $d(s)$ for the CKM-favored (-suppressed) $\Delta S=0\;(\Delta S=1)$ processes with $S$ being strange number.
@@
where the ratios $r_{2}=m_{L}/m_{B_{c}}$ and $r_{3}=m_{\eta_{c}}/m_{B_{c}}$ with $m_{L}(m_{\eta_{c}})$ being the light ($\eta_{c}$) meson mass, $P_{1}$ denotes the momentum carried by $B_{c}$ meson in its rest frame, and $P_{2}$ and $P_{3}$ denotes the momenta carried by $L$ and $\eta_{c}$ mesons moving along the plus and minus $z$ directions, respectively. Thanks to conservation of the angular momentum, the possible polarization vectors $\epsilon_{2L}$ ( here, the subscript $L$ stands for the longitudinal polarization, not to be confused with the abbreviation $L$ of light mesons) could be easily derived through the constraints $P_{2}\cdot\epsilon_{2L}=0$ and $\epsilon_{2L}^{2}=-1$. Notice that, if $L$ is a tensor meson, then a new longitudinal polarization vector could be constructed similarly relative to those of vectors and axial vectors but with an additional factor $\sqrt{2/3}$ Berger:2000wt ; Datta:2007yk ; Liu:2017cwl ; Liu:2023kxr . Then, specifically, $\epsilon_{2L}=\frac{1}{\sqrt{2(1-r_{3}^{2})}r_{2}}(1-r_{3}^{2},-r_{2}^{2},{\bf 0}_{T})$ for vectors and axial vectors, while $\epsilon_{2L}=\frac{1}{\sqrt{3(1-r_{3}^{2})}r_{2}}(1-r_{3}^{2},-r_{2}^{2},{\bf 0}_{T})$ for tensors in this work. The momenta of valence quarks in the initial- and final-state mesons are parametrized as
@@
The $B_{c}^{+}\to\eta_{c}L^{+}$ decay amplitude in the iPQCD formalism can therefore be conceptually written as,
@@
where $b_{i}$ is the conjugate space coordinate of transverse momentum $k_{iT}$, $t$ is the largest running energy scale in hard kernel $H(x_{i},b_{i},t)$, Tr denotes the trace over Dirac and $SU(3)$ color indices, $C(t)$ stands for the Wilson coefficients including the large logarithms $\ln(m_{W}/t)$ Keum:2000ph , and $\Phi$ is the wave function describing the hadronization of quark and antiquark to the meson. The Sudakov factor $e^{-S(t)}$ arises from $k_{T}$ resummation, which provides a strong suppression on the long distance contributions in the small $k_{T}$ (or large $b$) region Botts:1989kf . The detailed discussions for $e^{-S(t)}$ can be easily found in the original Refs. Botts:1989kf ; Liu:2018kuo ; Liu:2020upy . Thus, with Eq. (12), we can give the convoluted amplitudes of the decays $B_{c}^{+}\to\eta_{c}L^{+}$ explicitly through the evaluations of hard kernel $H(x_{i},b_{i},t)$ at leading order in the $\alpha_{s}$ expansion within the iPQCD formalism.
@@
where $P$ is the momentum of $B_{c}$ meson, $N_{c}=3$ is the color factor, and $x$ and ${\bf k}_{T}$ are the momentum fraction and intrinsic transverse momentum of charm quark in the $B_{c}$ meson. Here, $\phi_{B_{c}}(x,{\bf k}_{T})$ is the $B_{c}$-meson leading-twist distribution amplitude, whose explicit form in the impact $\bf b$} space is Liu:2018kuo
@@
with the decay constant $f_{B_{c}}=0.489\pm 0.005$ GeV Chiu:2007km , the shape parameter $\beta_{B_{c}}=1.0\pm 0.1$ GeV Liu:2018kuo , and $m_{c}$ and $m_{b}$ being the charm and bottom quark masses. Moreover, the normalization constant $N_{B_{c}}$ is fixed by the following relation:
@@
For the $\eta_{c}$ meson, its wave function has been studied within the nonrelativistic QCD approach Bondar:2004sv and derived as
@@
with $P$ and $m$ being the momentum and mass of $\eta_{c}$ and $x$ describing the charm-quark momentum fraction in $\phi^{v}_{\eta_{c}}(x)$ and $\phi^{s}_{\eta_{c}}(x)$ the twist-2 and twist-3 distribution amplitudes,
@@
The light-cone wave functions including distribution amplitudes for light pseudoscalars, scalars, vectors, axial vectors, and tensors calculated in the QCD sum rules up to twist 3 have been collected in Ref. Liu:2023kxr (and references therein). For simplicity, their explicit expressions would no longer be presented in this work. The readers could refer to Ref. Liu:2023kxr for detail.
Figure 1: Leading-order Feynman diagrams for the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism: (a) and (b) ((c) and (d)) contributing to the (non)factorizable-emission amplitudes.
The leading-order Feynman diagrams for the decays $B_{c}^{+}\to\eta_{c}L^{+}$ in the iPQCD formalism are shown in Fig. 1. Similar to those in Ref. Liu:2023kxr , we use $F_{e}$ and $M_{e}$ to describe the factorizable emission and the nonfactorizable emission amplitudes induced by the (VA)(VA)$(V-A)$ operators. The $B_{c}^{+}\to\eta_{c}L^{+}$ decay amplitude can thus be decomposed into
@@
in which, the related factorization formulas are given explicitly as follows:
  • (i)
@@
    where the ratios $r_{b}=m_{b}/m_{B_{c}}$ and $r_{c}=m_{c}/m_{B_{c}}$. The hard function $h_{i}(x_{i},b_{i})$ and the evolution function $E_{f}(t_{i})$ could refer to those expressions in Ref. Liu:2023kxr ,
@@
with the $B_{c}$-meson lifetime $\tau_{B_{c}}$ and the decay width $\Gamma$. Note that $|{\bf P}_{c}|\equiv|{\bf P}_{L}|=|{\bf P}_{\eta_{c}}|=\sqrt{\lambda(m_{B_{c}}^{2},m_{L}^{2},m_{\eta_{c}}^{2})}/(2m_{B_{c}})$ is the momentum of either $L$ or $\eta_{c}$ meson in the final states, with the $\rm K\ddot{a}ll\acute{e}n$} function $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$ ParticleDataGroup:2024cfk .
@@ III Numerical Results and Discussions
In numerical calculations, central values of the input parameters will be used implicitly unless otherwise stated. We adopt the relevant QCD scale (GeV), masses (GeV), and $B_{c}$-meson lifetime (ps) Keum:2000ph ; ParticleDataGroup:2024cfk ,
@@
In the following context, we will classify the considered decays into two different groups conventionally, that is, the decays $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ and the decays $B_{c}^{+}\to\eta_{c}(A,S,T)^{+}$, respectively, to present the related iPQCD predictions and phenomenological insights.
III.1 $B_{c}^{+}\to\eta_{c}(P,V)^{+}$
The CP-averaged $B_{c}^{+}\to\eta_{c}(\pi,K)^{+}$ BRs in the iPQCD formalism are presented as
@@ III.1 $B_{c}^{+}\to\eta_{c}(P,V)^{+}$
where the uncertainties are dominated by the shape parameter $\beta_{B_{c}}$ from the $B_{c}$-meson distribution amplitude. Note that, though having a smaller decay constant $f_{\eta_{c}}\sim 0.387$ than $f_{J/\psi}\sim 0.405$ and the same leading-twist distribution amplitudes in both $\eta_{c}$ and $J/\psi$, the $B_{c}^{+}\to\eta_{c}\pi^{+}$ BR is still larger than the $B_{c}^{+}\to J/\psi\pi^{+}$ one in the iPQCD formalism. In fact, from the numerical results of decay amplitudes shown in Table 1, it is found that ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ and ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})$ are governed by highly different QCD dynamics, that is, the former (latter) determined by the contributions arising from the distribution amplitude of twist-3 $\phi_{\eta_{c}}^{s}(x)$ [twist-2 $\phi_{J/\psi}^{L}(x)$] of $\eta_{c}(J/\psi)$, even with enhancement (reduction) by that of twist-2 $\phi_{\eta_{c}}^{v}(x)$ [twist-3 $\phi_{J/\psi}^{t}(x)$]. Any precise measurements in various experiments to help understand the QCD behavior of these $c\bar{c}$-mesons in depth are urgently demanded.
@@
According to the LHCb data LHCb:2014mvo , that is, $\frac{\sigma(B_{c}^{+})}{\sigma(B^{+})}\times{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})=(6.97\pm 0.13)\times 10^{-6}$ Liu:2025ipe in the fiducial region corresponding to the transverse momentum $0<p_{T}<20$ GeV and the rapidity $2.0<y<4.5$, and the assumption ${\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})\sim{\cal O}(10^{-3})$, then the forthcoming measurements of ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})$ in the LHCb experiment might be
@@
Based on the data for $\eta_{c}$ strong decays to stable hadrons ParticleDataGroup:2024cfk , one can derive the multibody $B_{c}$-meson decays via $\eta_{c}$ resonance under the narrow-width approximation:
@@
    lead to the three-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p})$ and five-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-})$ BRs as
@@
    ${\rm BR}(B_{c}^{+}\to\pi^{+}\eta_{c}(\to p\bar{p}\pi^{+}\pi^{-}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})\cdot{\cal B}(\eta_{c}\to p\bar{p}\pi^{+}\pi^{-})=(7.51^{+2.21}_{-1.83})\times 10^{-6}\;.$ (35)
@@
    result in the five-body $B_{c}^{+}\to\pi^{+}\eta_{c}(\to 2(\pi^{+}\pi^{-}),(\pi^{+}\pi^{-}K^{+}K^{-}),2(K^{+}K^{-}))$ BRs as
@@
    $=$ $(2.84^{+1.10}_{-0.99})\times 10^{-6}\;.$
In principle, the related measurements are previously difficult due to the large background and the small efficiency at LHC. However, these BRs around ${\cal O}(10^{-6})$ and above with fully charged final states are expected to be probed in the near future, since the LHCb detector has accomplished a successful upgrade in 2022.
@@
The $B_{c}^{+}\to\eta_{c}V^{+}$ BRs in the iPQCD formalism can be read as
@@
where the theoretical errors are also dominated mainly by the $B_{c}$-meson shape parameter $\beta_{B_{c}}$. By employing Eqs. (33) and (36), the measurable BRs of multibody $B_{c}^{+}\to\eta_{c}\rho^{+}$ decays through resonance state $\eta_{c}$ could be derived under the narrow-width approximation as follows:
@@
The ratio between the $B_{c}^{+}\to\eta_{c}\rho^{+}$ and $B_{c}^{+}\to(J/\psi,\eta_{c})\pi^{+}$ BRs could be deduced as
@@
Here, the latter ratio $R_{\rho/\pi}^{\eta_{c}}$ is close to that obtained from the $B_{c}^{+}\to J/\psi(\pi,\rho)^{+}$ decays Liu:2023kxr . Moreover, the ratio between the $B_{c}^{+}\to\eta_{c}K^{*+}$ and $B_{c}^{+}\to\eta_{c}\rho^{+}$ BRs is